Joe's towing company charges a base rate of $90 plus $4.50 per mile for each mile over 2 miles. Mac"s towing company charges a base rate of $70 plus $5 per mile for every mile over 2 miles. For what total mileage will both companies charge the same amount? (set up a SYSTEM OF EQUATIONS and solve)
Answer by richwmiller(17219)

4.5x+90=5x+70
20=0.5x
20/0.5=x
x=40.0
y=4.5*40+90
y=5*40+70
y=270
They are equal when there are 42 miles x=40 +2miles
They are equal when total cost y=$270